OpenStudy (anonymous):

2(x+4)^2=9 I need to solve using the square root procedure....looking for help in how to work this

OpenStudy (whpalmer4):

\[2(x+4)^2 = 9\]Divide both sides by 2\[(x+4)^2=\frac{9}{2}\]Now take the square root of both sides \[(x+4) = \pm\sqrt{\frac{9}{2}}\]Can you do the rest?

OpenStudy (anonymous):

That's how I had it done but the answer has a -8 in it and I couldn't figure out where it came from

OpenStudy (whpalmer4):

Okay. Let's go through it. \[x+4 = \pm\sqrt{\frac{9}{2}}\]\[x = -4\pm\frac{3}{\sqrt{2}}\]Rationalize the denominator by multiplying by \(\dfrac{\sqrt{2}}{\sqrt{2}}\) \[x = -4 \pm \frac{3\sqrt{2}}{2} = \frac{-8\pm 3\sqrt{2}}{2}\]

OpenStudy (whpalmer4):

Now does it look the way you expect?

OpenStudy (anonymous):

that -8 is still throwing me. I understand -4x2= the -8. but I'm not sure if I saw another problem like this, that I would know what to multiply to get that last part

OpenStudy (anonymous):

thank you so much for the help though, it's more clear than it was before :)

OpenStudy (whpalmer4):

yeah, a lot of this is just comfort manipulating fractions and radicals...

OpenStudy (whpalmer4):

not uncommon to end up with the right answer, just in a slightly different form that looks very dissimilar :-)

OpenStudy (whpalmer4):

one thing you can do, assuming your calculator skills are good, is evaluate the expression with a calculator and see if it works out to be the same number.

OpenStudy (whpalmer4):

gotta run, good luck!

OpenStudy (anonymous):

Thank you!